An athlete executing a long jump leaves the ground at a 28.0 angle and travels 7.80 m. (a) What was the takeoff speed? (b) If this speed were increased just 50%, how much longer would the jump be?
1
Expert's answer
2021-10-24T18:26:33-0400

Given:

"\theta=28.0^{\circ}"

"R=7.80\:\rm m"


(a) The range

"R=\frac{v_0^2\sin2\theta}{g}"

So, the initial speed

"v_0=\sqrt{\frac{gR}{\sin2\theta}}=\sqrt{\frac{9.8*7.8}{\sin2*28^{\circ}}}=9.6\:\rm m/s"

(b) The new range

"R_1=\frac{(1.5v_0)^2\sin2\theta}{g}=1.5^2R=1.5^2*7.8=17.6\:\rm m"